OpenStudy (anonymous):

How do i find the derivative of (6x / the square root of x squared + 5)?

OpenStudy (anonymous):

Quotient rule: Assuming \[f(x)=\frac{6x}{\sqrt{x^2+5}}\] Then: \[f'(x)=\frac{6(\sqrt{x^2+5})-(6x)(\frac{1}{2})(2x)(x^2+5)^{\frac{-1}{2}}}{(\sqrt{x^2+5})^2}\] Then you can simplify it from there.

OpenStudy (anonymous):

Refresh if the latex looks bad

OpenStudy (anonymous):

this site is very helpful: http://www.mathwords.com/d/derivative_rules.htm

OpenStudy (amistre64):

orrr, go with the whole f(x+h)-f(x) fiasco :)

OpenStudy (anonymous):

I agree with amistre^^

OpenStudy (anonymous):

Builds character :)

OpenStudy (amistre64):

if you cant remember the shortcuts; you pretty much dont deserve them ;)

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

Thanx